OpenStudy (anonymous):

Consider the following predator-prey system where x and y are in millions of creatures and t represents time in years: dy/dt=2x-xy dy/dt=-4y+xy 1. Show that (4,2) is the nonzero equilibrium solution . 2. Find an expression for dy/dx

OpenStudy (anonymous):

The first must be dx/dt!

OpenStudy (anonymous):

That is what I confused too-.-

OpenStudy (anonymous):

Do you know how to solve?

OpenStudy (anonymous):

Let me have try

OpenStudy (anonymous):

I need the second part Thank you

OpenStudy (anonymous):

it must be wrong!you can see it: http://en.wikipedia.org/wiki/Lotka%E2%80%93Volterra_equation

OpenStudy (anonymous):

Which one is the real \(\frac{dx}{dt}\)?

OpenStudy (anonymous):

Also, \[\frac{dy}{dx} = \dfrac{\left(\dfrac{dy}{dt}\right)}{\left(\dfrac{dx}{dt}\right)}\]